a jar containing only nickels and dimes contains 60 coins. The value of all the coins in the jar is $4.45. solve by elimination to find the amount of nickels and dimes are in the jar

Answer by checkley77(12844)

N+D=60 MULTIPLY THIS EQUATION BY -.05 & ADD
.05N+.10D=4.45
-.05N-.05D=-3
-----------------
.05D=1.45
D=1.45/.05
D=29 DIMES.
60-29=31 NICKELS.
PROOF:
.05*31+.10*29=4.45
1.55+2.90=4.45
4.45=4.45